# Determination of the stably free cancellation property for orders

### v2

@@ 9.1. An algorithm to show that an order fails SFC
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. If the following algorithm returns $\mathsf{false}$, then $\Lambda$ fails SFC. If it returns $\mathsf{true}$, then $\Lambda$ is maximal and has SFC. If it returns $\mathsf{inconclusive}$, then nothing can be said about whether $\Lambda$ has SFC or not.
@@
In step (2), the correctness of an output $\mathsf{false}$ follows from Corollary 3.2 and the correctness of an output $\mathsf{true}$ is clear. Step (4) ensures that $X_{\beta_{i}}$ is stably free over $\Lambda$ for each $i\in\{1,\ldots,r\}$ by Proposition 8.3 (v),(vii). Thus output of step (5) is correct because if there exists $i\in\{1,\ldots,r\}$ such that $X_{\beta_{i}}$ is not free over $\Lambda$, then $\Lambda$ clearly fails SFC. Note that step (2) is in fact optional, in that the algorithm is still correct even if it is omitted. ∎
@@ 9.2. Integral group rings without SFC
In order, the groups listed in Theorem 9.4 (i) are isomorphic to $G_{(16,12)}$, $G_{(24,12)}$, $G_{(32,41)}$, $G_{(40,7)}$, and those in (ii) to $G_{(96,198)}$, $G_{(96,188)}$, $G_{(480,960)}$.
@@
We first show that $\mathbb{Z}[G]$ fails SFC. Let $e$ be the central idempotent of $\mathbb{Q}[G]$ such that $e\mathbb{Q}[G]$ is the direct sum of all simple components of $\mathbb{Q}[G]$ that are either commutative or totally definite quaternion algebras. For each group $G$ listed apart from $G=\tilde{O}\times C_{2}$, verification that $e\mathbb{Z}[G]$ fails SFC was performed using Algorithm 9.1. Theorem 3.1 then implies that $\mathbb{Z}[G]$ also fails SFC. For $G=\tilde{O}\times C_{2}$ there exists a normal subgroup $N$ of order $2$ such that $G/N\cong S_{4}\times C_{2}$. For this choice of $G$ and $N$, the verification that the order $\Gamma:=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$ fails SFC was performed using Algorithm 9.1; again, Theorem 3.1 then implies that $\mathbb{Z}[G]$ also fails SFC.
@@
$Q_{8},\,Q_{12},\,Q_{16},\,Q_{20},\,\tilde{T},\,\tilde{O},\,\tilde{I},$ (9.1)
@@
10. Determining SFC using fibre products
@@ 10.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. Let $I$ and $J$ be two-sided ideals of $\Lambda$ such that $I\cap J=0$ and $K(I+J)=A$. Let $\Lambda_{1}=\Lambda/I$, let $\Lambda_{2}=\Lambda/J$ and let $\overline{\Lambda}=\Lambda/(I+J)$. Then $\Lambda_{1}$ and $\Lambda_{2}$ are $\mathcal{O}$-orders and $\overline{\Lambda}$ is an $\mathcal{O}$-torsion $\mathcal{O}$-algebra. Then as in Example 5.1, we have a fibre product
${\Lambda}$${\Lambda_{1}}$${\Lambda_{2}}$${\overline{\Lambda},}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$ (10.1)
@@
Let $A_{1}:=K\Lambda_{1}\cong KJ$ and let $A_{2}:=K\Lambda_{1}\cong KI$. Then by making the appropriate identifications, we can and do assume that
$\Lambda=\{(x_{1},x_{2})\in\Lambda_{1}\oplus\Lambda_{2}\mid g_{1}(x_{1})=g_{2}(x_{2})\}\subseteq\Lambda_{1}\oplus\Lambda_{2}\subseteq A_{1}\oplus A_{2}=A.$ (10.2)
@@
We shall relate pullback lattices to test lattices and use this to give an alternative description of $\ker(\partial)$. Under the further assumption that $A_{1}$ satisfies the Eichler condition, this will lead to an algorithm to test whether $\Lambda$ has SFC or not.
@@ 10.2. Test lattices and pullback lattices
For $i=1,2$, let $\mathcal{M}_{i}$ be a choice of $\mathcal{O}$-order in $A_{i}$ containing $\Lambda_{i}$ (such choices exist by [Rei03, (10.4)]) and let $\mathfrak{f}_{i}$ be any proper full two-sided ideal of $\mathcal{M}_{i}$ that is contained in $\Lambda\cap A_{i}$. Let $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$ and let $\mathfrak{f}=\mathfrak{f}_{1}\oplus\mathfrak{f}_{2}$. For $i=1,2$, let $C_{i}$ be the centre of $A_{i}$ and let $\mathfrak{g}_{i}:=\mathfrak{f}_{i}\cap\mathcal{O}_{C_{i}}=\mathfrak{f}_{i}\cap C_{i}$. Let $C=C_{1}\oplus C_{2}$ and let $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$. We set
@@
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\beta=(\beta_{1},\beta_{2})\in\Lambda_{1}\oplus\Lambda_{2}$ with $\beta\in A^{\times}$ and $\overline{\beta_{i}}\in\left(\Lambda_{i}/\mathfrak{f}_{i}\right)^{\times}$ for $i=1,2$. Let $u=g_{1}(\beta_{1})g_{2}(\beta_{2})^{-1}\in\overline{\Lambda}^{\times}$. Then $X_{\beta}=M(u)\beta$, so in particular, $X_{\beta}\cong M(u)$ as $\Lambda$-lattices.
@@
Suppose that $\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$. Since $\Lambda_{\mathfrak{p}}\subseteq\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}$ and $\Lambda_{\mathfrak{p}}$ is maximal, we have that $\Lambda_{\mathfrak{p}}=\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}$. Thus localising the fibre product (5.1) at $\mathfrak{p}$ implies that $\mathfrak{h}_{i,\mathfrak{p}}=\mathcal{M}_{i,\mathfrak{p}}=\Lambda_{i,\mathfrak{p}}$ for $i=1,2$ and $\overline{\Lambda}_{\mathfrak{p}}=0$. It is then easy to see that $M(u)_{\mathfrak{p}}=\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$ and hence $M(u)_{\mathfrak{p}}\beta=\mathcal{M}_{\mathfrak{p}}\beta=(X_{\beta})_{\mathfrak{p}}$, where the last equality follows from Lemma 8.2.
@@
$M(u)_{\mathfrak{p}}\beta$ =$=\{(a_{1}\beta_{1},a_{2}\beta_{2})\mid(a_{1},a_{2})\in\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}\text{ and }g_{1,\mathfrak{p}}(a_{1})u=g_{2,\mathfrak{p}}(a_{2})\}$
=$=\{(a_{1}\beta_{1},a_{2}\beta_{2})\mid(a_{1},a_{2})\in\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}\text{ and }g_{1,\mathfrak{p}}(a_{1}\beta_{1})=g_{2,\mathfrak{p}}(a_{2}\beta_{2})\}$
=$=\{(b_{1},b_{2})\mid(b_{1},b_{2})\in\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}\text{ and }g_{1,\mathfrak{p}}(b_{1})=g_{2,\mathfrak{p}}(b_{2})\}$
@@ 10.3. Alternative description of $\ker(\partial)$
$\nu_{2}^{\prime}\colon(\Lambda_{2}/\mathfrak{h}_{2})^{\times}\longrightarrow(I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+})/\tilde{\nu}((\Lambda/\mathfrak{f})^{\times}),\quad x_{2}\mapsto\tilde{\mu}(\tilde{\nu}((\overline{1},\overline{\alpha_{2}}))),$ (10.3)
@@
The map $\nu_{2}^{\prime}$ is a well-defined group homomorphism. If $\mathfrak{g}$ satisfies Assumption (G) then $\ker(\partial)={\tilde{g}_{2}}(\ker(\nu^{\prime}_{2}))$, where $\tilde{g}_{2}:\Lambda_{2}/\mathfrak{h}_{2}\cong\overline{\Lambda}$ is the ring isomorphism induced by $g_{2}$.
@@
10.4. An algorithm for determining SFC using fibre products
@@
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$ with a fibre product as in (10.1). Assume that $A_{1}$ is nontrivial and satisfies the Eichler condition. If the following algorithm returns $\mathsf{true}$, then $\Lambda$ has SFC. If it returns $\mathsf{false}$, then $\Lambda$ fails SFC.
@@
    Compute the centre $C=C_{1}\oplus C_{2}$ of $A=A_{1}\oplus A_{2}$ and the decomposition of $A=A_{1}\oplus A_{2}$ into simple $K$-algebras.
@@
    For $i=1,2$, compute $\mathfrak{f}_{i}:=\{x\in A_{i}\mid x\mathcal{M}_{i}\subseteq\Lambda\}\cap\{x\in A_{i}\mid\mathcal{M}_{i}x\subseteq\Lambda_{i}\}$.
@@
    For $i=1,2$, compute $\mathcal{O}_{C_{i}}:=\mathcal{M}_{2}\cap C_{i}$ and $\mathfrak{g}_{i}:=\mathfrak{f}_{i}\cap\mathcal{O}_{C_{i}}=\mathfrak{f}_{i}\cap C_{i}$.
@@
    For $i=1,2$, adjust $\mathfrak{f}_{i}$ (and thus $\mathfrak{g}_{i}$) to ensure that $\mathfrak{g}_{i}$ satisfies Assumption (G).
@@
In principle, Step (7) can be performed using any appropriate algorithm, such as Algorithm 8.9, for example. Note that Algorithm 10.3 itself can be applied recursively, which is useful because the choice of fibre product can have a large impact on the running time.
@@ 11.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$ and let $\mathfrak{h}$ be any proper full two-sided ideal of $\Lambda$. In §11.2, we describe a practical algorithm to compute a set of generators of the unit group $(\Lambda/\mathfrak{h})^{\times}$. In §11.3, we then describe a practical algorithm to compute $(\Lambda/\mathfrak{h})^{\times}$ as an abstract group. These are important subtasks in Algorithm 10.3.

### v3

@@ 9.1. An algorithm to show that an order fails SFC
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. If the following algorithm returns $\mathsf{false}$, then $\Lambda$ fails SFC. If it returns $\mathsf{true}$, then $\Lambda$ is maximal and has SFC. If it returns $\mathsf{inconclusive}$, then nothing can be said about whether $\Lambda$ has SFC.
@@
In step (2), the correctness of an output $\mathsf{false}$ follows from Corollary 3.2 and the correctness of an output $\mathsf{true}$ is clear. Step (4) ensures that $X_{\beta_{i}}$ is stably free over $\Lambda$ for each $i\in\{1,\ldots,r\}$ by Proposition 8.3 (v),(vii). Thus the output of step (5) is correct because if there exists $i\in\{1,\ldots,r\}$ such that $X_{\beta_{i}}$ is not free over $\Lambda$, then $\Lambda$ clearly fails SFC. Note that step (2) is in fact optional, in that the algorithm is still correct even if it is omitted. ∎
@@ 9.2. Integral group rings without SFC
In order, the groups listed in Theorem 9.4 (i) are isomorphic to $G_{(16,12)}$, $G_{(24,7)}$, $G_{(32,41)}$, $G_{(40,7)}$, and those in (ii) to $G_{(96,198)}$, $G_{(96,188)}$, $G_{(480,960)}$.
@@
We first show that $\mathbb{Z}[G]$ fails SFC. Further details on the computations used in this part of the proof can be found in Appendix A. Let $e$ be the central idempotent of $\mathbb{Q}[G]$ such that $e\mathbb{Q}[G]$ is the direct sum of all simple components of $\mathbb{Q}[G]$ that are either commutative or totally definite quaternion algebras. For each group $G$ listed apart from $G=\tilde{O}\times C_{2}$, verification that $e\mathbb{Z}[G]$ fails SFC was performed using Algorithm 9.1. Theorem 3.1 then implies that $\mathbb{Z}[G]$ also fails SFC. For $G=\tilde{O}\times C_{2}$ there exists a normal subgroup $N$ of order $2$ such that $G/N\cong S_{4}\times C_{2}$. For this choice of $G$ and $N$, the verification that the order $\Gamma:=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$ fails SFC was performed using Algorithm 9.1; again, Theorem 3.1 then implies that $\mathbb{Z}[G]$ also fails SFC.
@@
(9.1) $Q_{8},\,Q_{12},\,Q_{16},\,Q_{20},\,\tilde{T},\,\tilde{O},\,\tilde{I}$,
@@
10. Determining SFC using fiber products
@@ 10.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. Let $I$ and $J$ be two-sided ideals of $\Lambda$ such that $I\cap J=0$ and $K(I+J)=A$. Let $\Lambda_{1}=\Lambda/I$, let $\Lambda_{2}=\Lambda/J$ and let $\overline{\Lambda}=\Lambda/(I+J)$. Then $\Lambda_{1}$ and $\Lambda_{2}$ are $\mathcal{O}$-orders and $\overline{\Lambda}$ is an $\mathcal{O}$-torsion $\mathcal{O}$-algebra. Then as in Example 5.1, we have a fiber product
(10.1) $\Lambda$$\Lambda_{1}$$\Lambda_{2}$$\overline{\Lambda},$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$
@@
Let $A_{1}:=K\Lambda_{1}\cong KJ$ and let $A_{2}:=K\Lambda_{2}\cong KI$. Then by making the appropriate identifications, we can and do assume that
(10.2) $\Lambda=\{(x_{1},x_{2})\in\Lambda_{1}\oplus\Lambda_{2}\mid g_{1}(x_{1})=g_{2}(x_{2})\}\subseteq\Lambda_{1}\oplus\Lambda_{2}\subseteq A_{1}\oplus A_{2}=A.$
@@
We shall relate pullback lattices to test lattices and use this to give an alternative description of $\ker(\partial)$. Under the further assumption that $A_{1}$ satisfies the Eichler condition, this will lead to an algorithm to test whether $\Lambda$ has SFC.
@@ 10.2. Test lattices and pullback lattices
For $i=1,2$, let $\mathcal{M}_{i}$ be a choice of maximal $\mathcal{O}$-order in $A_{i}$ containing $\Lambda_{i}$ (such choices exist by [Rei03, (10.4)]) and let $\mathfrak{f}_{i}$ be any full two-sided ideal of $\mathcal{M}_{i}$ that is contained in $\Lambda\cap A_{i}$. Let $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$ and let $\mathfrak{f}=\mathfrak{f}_{1}\oplus\mathfrak{f}_{2}$. For $i=1,2$, let $C_{i}$ be the center of $A_{i}$ and let $\mathfrak{g}_{i}:=\mathfrak{f}_{i}\cap\mathcal{O}_{C_{i}}=\mathfrak{f}_{i}\cap C_{i}$. Let $C=C_{1}\oplus C_{2}$ and let $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$. We set
@@
Suppose that $\mathfrak{g}$ is locally radical. Let $\beta=(\beta_{1},\beta_{2})\in\Lambda_{1}\oplus\Lambda_{2}$ with $\beta\in A^{\times}$ and $\overline{\beta_{i}}\in\left(\Lambda_{i}/\mathfrak{f}_{i}\right)^{\times}$ for $i=1,2$. Let $u=g_{1}(\beta_{1})g_{2}(\beta_{2})^{-1}\in\overline{\Lambda}^{\times}$. Then $X_{\beta}=M(u)\beta$, so in particular, $X_{\beta}\cong M(u)$ as $\Lambda$-lattices.
@@
Suppose that $\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$. Since $\Lambda_{\mathfrak{p}}\subseteq\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}$ and $\Lambda_{\mathfrak{p}}$ is maximal, we have that $\Lambda_{\mathfrak{p}}=\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}$. Thus localising the fiber product (5.1) at $\mathfrak{p}$ implies that $\mathfrak{h}_{i,\mathfrak{p}}=\mathcal{M}_{i,\mathfrak{p}}=\Lambda_{i,\mathfrak{p}}$ for $i=1,2$ and $\overline{\Lambda}_{\mathfrak{p}}=0$. It is then easy to see that $M(u)_{\mathfrak{p}}=\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$ and hence $M(u)_{\mathfrak{p}}\beta=\mathcal{M}_{\mathfrak{p}}\beta=(X_{\beta})_{\mathfrak{p}}$, where the last equality follows from Lemma 8.2.
@@
$M(u)_{\mathfrak{p}}\beta$ $=\{(a_{1}\beta_{1},a_{2}\beta_{2})\mid(a_{1},a_{2})\in\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}\text{ and }g_{1,\mathfrak{p}}(a_{1})u=g_{2,\mathfrak{p}}(a_{2})\}$
$=\{(a_{1}\beta_{1},a_{2}\beta_{2})\mid(a_{1},a_{2})\in\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}\text{ and }g_{1,\mathfrak{p}}(a_{1}\beta_{1})=g_{2,\mathfrak{p}}(a_{2}\beta_{2})\}$
$=\{(b_{1},b_{2})\mid(b_{1},b_{2})\in\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}\text{ and }g_{1,\mathfrak{p}}(b_{1})=g_{2,\mathfrak{p}}(b_{2})\}$
@@ 10.3. Alternative description of $\ker(\partial)$
(10.3) $\nu_{2}^{\prime}\colon(\Lambda_{2}/\mathfrak{h}_{2})^{\times}\longrightarrow(I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+})/\tilde{\nu}((\Lambda/\mathfrak{f})^{\times}),\quad x_{2}\mapsto\tilde{\mu}(\tilde{\nu}((\overline{1},\overline{\alpha_{2}}))),$
@@
The map $\nu_{2}^{\prime}$ is a well-defined group homomorphism. If $\mathfrak{g}$ is locally radical then $\ker(\partial)={\tilde{g}_{2}}(\ker(\nu^{\prime}_{2}))$, where $\tilde{g}_{2}:\Lambda_{2}/\mathfrak{h}_{2}\cong\overline{\Lambda}$ is the ring isomorphism induced by $g_{2}$.
@@
10.4. An algorithm for determining SFC using fiber products
@@
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$ with a fiber product as in (10.1). Assume that $A_{1}$ is nontrivial and satisfies the Eichler condition. If the following algorithm returns $\mathsf{true}$, then $\Lambda$ has SFC. If it returns $\mathsf{false}$, then $\Lambda$ fails SFC.
@@
    Compute the center $C=C_{1}\oplus C_{2}$ of $A=A_{1}\oplus A_{2}$ and the decomposition of $A=A_{1}\oplus A_{2}$ into simple $K$-algebras.
@@
    For $i=1,2$, compute $\mathfrak{f}_{i}:=\{x\in A_{i}\mid x\mathcal{M}_{i}\subseteq\Lambda_{i}\}\cap\{x\in A_{i}\mid\mathcal{M}_{i}x\subseteq\Lambda_{i}\}$.
@@
    For $i=1,2$, compute $\mathcal{O}_{C_{i}}:=\mathcal{M}_{i}\cap C_{i}$ and $\mathfrak{g}_{i}:=\mathfrak{f}_{i}\cap\mathcal{O}_{C_{i}}=\mathfrak{f}_{i}\cap C_{i}$.
@@
    For $i=1,2$, adjust $\mathfrak{f}_{i}$ (and thus $\mathfrak{g}_{i}$) to ensure that $\mathfrak{g}_{i}$ is locally radical.
@@
In principle, Step (7) can be performed using any appropriate algorithm, such as Algorithm 8.9, for example. Note that Algorithm 10.3 itself can be applied recursively, which is useful because the choice of fiber product can have a large impact on the running time.
@@ 11.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$ and let $\mathfrak{h}$ be any full two-sided ideal of $\Lambda$. In §11.2, we describe a practical algorithm to compute a set of generators of the unit group $(\Lambda/\mathfrak{h})^{\times}$. In §11.3, we then describe a practical algorithm to compute $(\Lambda/\mathfrak{h})^{\times}$ as an abstract group. These are important subtasks in Algorithm 10.3.